# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v2

@@ Abstract
Molecules that violate Hund’s rule—having first excited singlet state (S1) below the triplet state (T1)—are rare yet promising as efficient light emitters. Their high-throughput identification demands exceptionally accurate excited-state modeling to minimize false positives and negatives. Benchmarking twelve S1–T1 energy gaps, we find that local variants of ADC(2) and CC2 deliver excellent accuracy and speed for screening medium-sized molecules. Notably, while double-hybrid DFT approximations (e.g., B2GP-PLYP and PBE-QIDH) exhibit high mean errors ($>100$ meV) despite very low standard deviations ($\approx 10$ meV), exploring their parameter space reveals that a configuration with 75% exchange and 55% correlation reduces the mean error to below $5$ meV—albeit with increased variance. Using this low-bias parameterization as an internal reference, we correct the systematic error while maintaining low variance, effectively combining the strengths of both low-bias and low-variance DFT parameterizations to enhance overall accuracy. Our findings suggest that low-variance DFT methods—often overlooked due to high bias—can serve as reliable tools for predictive modeling in first-principles molecular design.
1 Introduction
@@
The energy criterion for organic light-emitting diodes (OLEDs) to operate via thermally activated delayed fluorescence (TADF) is that the energy of the first excited singlet state (S1) must be no more than $0.1$ eV higher than that of the triplet state (T1)[1, 2]. This condition facilitates thermally-aided reverse intersystem crossing (RISC), enabling the theoretical complete transfer of the T1 population to the emissive state, S1[3]. A derivative of the triangular molecule, heptazine (aka heptaazaphenalene, 7AP), with a central nitrogen (N) atom in an anti-aromatic 12-annulene framework, showed delayed fluorescence without thermal assistance, indicating a negative S1–T1 gap (STG)[4]. This mechanism paves the way for designing OLEDs that leverage exothermic delayed fluorescence from inverted singlet and triplet excited states (DFIST). Direct spectroscopic evidence of a negative STG ($-0.047\pm 0.007$ eV) was obtained for pentaazaphenalene (5AP) using anion photoelectron spectroscopy and fluorescence measurements [5]. Negative STG was also was also observed in the transient photoluminescence spectrum of dialkylamine-substituted 5AP[6]. Computational studies have further suggested that azaphenalenes (APs) with other substitution patterns, as well as the boron (B) analog of azaphenalene (1AP), boraphenalene (1BP), also exhibit negative STGs[7, 8, 9, 10, 11, 12, 13, 14]. Non-alternant hydrocarbons and their substituted analogs show the potential for negative STGs[15, 16, 17, 18, 19]. Notably, substituted analogs of the non-fused bicyclic hydrocarbon have demonstrated negative STGs, attributed to through-bond charge-transfer states[20].
@@
A comprehensive search across $12,880$ structurally diverse small molecules revealed no exceptions to Hund’s rule, indicating that achieving the electronic structure criteria for ${\rm STG}<0$ requires nontrivial molecular structures[21]. Moreover, the typical magnitudes of negative STGs are much smaller or similar to the average errors associated with popular quantum chemistry and density functional theory approximations (DFAs). Consequently, due to the rarity of molecules with negative STGs and the limitations of current quantum chemical approximations, accurately identifying molecules exhibiting ${\rm STG}<0$ via high-throughput screening—with minimal false-positive and false-negative predictions—remains challenging. The linear-response time-dependent density functional theory (LR-TD-DFT) formalism provides qualitative accuracy—i.e., gets the sign of the STG for DFIST candidates correct—only within the double-hybrid (dh) DFT framework containing many-body correlation via the second-order perturbation theory (MP2). With the right combination of ingredients, dh-DFT can reliably predict a wide range of molecular properties[22, 23, 24, 23]. Nevertheless, selecting suitable exchange-correlation (XC) functionals within the dh-DFT framework remains challenging, especially for novel problems lacking sufficient reference data.
@@
We present a scheme that leverages the bias-variance tradeoff in the dh-DFT formalism. We evaluate various computational methods for predicting twelve STGs in triangular molecules, benchmarking them against previously reported theoretical best estimates[14]. We examine dh-DFAs exhibiting very low variance for STG predictions, and reparametrize them to minimize the bias. These optimal low-bias models are used as internal references to correct the predictions of the low-variance models. This internally referenced scaling approach is demonstrated by predicting STGs with low bias and low variance, resulting in excellent overall accuracy. We discuss the merits and drawbacks of this approach for its further applications.
2 Dataset and Computational Methods
We consider twelve systems (see Figure 1) using equilibrium geometries from Ref. [14], determined at the CCSD(T) level with the frozen core approximation and the cc-pVTZ basis set. Using these geometries, we performed single-point excited state calculations with various methods. In addition, we collected TBE of the S1 and T1 energies from Ref. [14] to serve as reference values for benchmarking. As stated in Ref. [14], structure 11 in Figure 1 (with $D_{3\rm h}$ symmetry) corresponds to a saddle point on the potential energy surface at the MP2/6-311G($d$,$p$) level, whereas the minimum energy geometry exhibits $C_{3\rm h}$ symmetry (structure 12). Similarly, a separate study[25] found that the high-symmetry forms of structures 6, 7, and 8 are not true minima and structure 2 is a very shallow potential well at the CCSD(T)/cc-pVTZ level[25]. Overall, four of the twelve structures are transition states; however, we include them in our benchmarking of STGs due to the availability of TBE-level results and the lack of alternative high-quality data.
@@ 2 Dataset and Computational Methods
Figure 1: Nitrogen- and boron-centered [12]-annulene systems reported in Ref. [14].
@@
Laplace-transformed, density-fitted, local versions of correlated methods—CC2 [26] and ADC(2) [27]—offer attractive speedups for modeling excited states of medium-sized molecules. We denote these methods as L-CC2 and L-ADC(2), and performed calculations with the cc-pVDZ, aug-cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets, using the corresponding JKFIT and MP2FIT auxiliary basis sets. L-CC2 and L-ADC(2) are the same methods denoted as LT-DF-CC2 and LT-DF-ADC(2) elsewhere. L-CC2 and L-ADC(2) calculations were conducted using Molpro (version 2015.1) [28]. L-CC2 has been shown to deviate from the canonical variant CC2 by $\leq 0.05$ eV, irrespective of the character of the excited states, including that of charge-transfer states for which CC2 can have huge error [29]. The systems studied in this work do not have charge-transfer characteristics, hence CC2 delivers good accuracy for STGs in agreement with TBEs[14]. L-CC2 has demonstrated excellent accuracy in modeling the excited states of BODIPY derivatives [30, 31]. For comparison, we collected S1 and T1 energies obtained using the linear-response coupled-cluster methods, EOM-CCSD and CC3, with the aug-cc-pVDZ basis set from Ref. [14]. Additionally, we performed ADC(2) calculations using the cc-pVDZ, cc-pVTZ, aug-cc-pVDZ, and aug-cc-pVTZ basis sets with the resolution-of-the-identity (RI) approximation [32, 33] as implemented in QChem (version 6.0.2) [34].
@@
We conducted LR-TD-DFT calculations using the Tamm–Dancoff approximation (TDA). Although TDA may not predict oscillator strengths as accurately as the random-phase approximation (RPA), RPA predictions of triplet energies are sensitive to triplet stability [35]. Therefore, TDA is preferred for automated high-throughput calculations, and previous work has shown that its STG predictions for azaphenalenes are more accurate than that of RPA[21]. The performance of various DFAs for predicting STGs in azaphenalines has been discussed in prior studies [36, 37, 38].
@@
This study explores several DFAs spanning all levels of the Jacob’s ladder [39], namely, local density approximation (LDA): VWN5 [40] and PWLDA; generalized gradient approximation (GGA): BP86, BLYP, PW91 [41, 42], PBE [43], mPWPW, and mPWLYP; hybrid GGA: B3LYP [44], PBE0 [45], X3LYP [46], mPW1PW, mPW1LYP, and BHandHLYP [47]; meta-GGA: TPSS [48], TPSSh [48], TPSS0 [48], and M06L [49] hybrid meta-GGA: M062X [50]; range-separated hybrids: $\omega$B97, [51] $\omega$B97X, [51]$\omega$B97X-D3, CAM-B3LYP [52], LC-BLYP [53], and LC-PBE [54]; double hybrid: $\omega$B97X-2 [55], B2PLYP [56], mPW2PLYP [57], B2GP-PLYP [58], PWPB95 [23], PBE-QIDH [59], PBE0-DH [60], $\omega$B2PLYP [61], $\omega$B2GP-PLYP [61], $\omega$B88PP86 [62], $\omega$PBEPP86 [62], SCS/SOS-B2PLYP21 [62], SCS-PBE-QIDH [62], SCS-B2GP-PLYP21 [62], SOS-B2GP-PLYP21 [62], SCS-$\omega$B2GP-PLYP21 [62], SOS-$\omega$B2GP-PLYP21 [62], SCS-RSX-QIDH [62], SOS-RSX-QIDH [62], SCS-$\omega$B88PP86 [62], SOS-$\omega$B88PP86 [62], SCS-$\omega$PBEPP86 [62], SOS-$\omega$PBEPP86 [62], DSD-BLYP [63], DSD-PBEP86 [63], DSD-PBEB95 [63], RSX-QIDH [64], and RSX-0H [65]. All DFT calculations were performed using Orca (version 6.0.0) [66, 67].
3 Bias Correction for Low-Variance Predictors
We examine the conditions under which regression-based bias correction is effective, to correct a low-variance, high-bias predictor using a low-bias, high-variance predictor. We clarify when and how the transformation preserves variance while ensuring bias reduction.
Definitions:

### v3

@@ Abstract
Molecules that have been suggested to violate the Hund’s rule, having a first excited singlet state (S1) energetically below the triplet state (T1), are rare. Yet, they hold the promise to be efficient light emitters. Their high-throughput identification demands exceptionally accurate excited-state modeling to minimize qualitatively wrong predictions. We benchmark twelve S1–T1 energy gaps to find that the local-correlated versions of ADC(2) and CC2 excited state methods deliver excellent accuracy and speed for screening medium-sized molecules. Notably, we find that double-hybrid DFT approximations (e.g., B2GP-PLYP and PBE-QIDH) exhibit high mean absolute errors ($>100$ meV) despite very low standard deviations ($\approx 10$ meV). Exploring their parameter space reveals that a configuration with 75% exchange and 55% correlation, which reduces the mean absolute error to below $5$ meV, but with an increased variance. Using this low-bias parameterization as an internal reference, we correct the systematic error while maintaining low variance, effectively combining the strengths of both low-bias and low-variance DFT parameterizations to enhance overall accuracy. Our findings suggest that low-variance DFT methods, often overlooked due to their high bias, can serve as reliable tools for predictive modeling in first-principles molecular design. The bias-correction data-fitting procedure can be applied to any general problem where two flavors of a method, one with low bias and another with low variance, have been identified a priori.
Electronic excited states, Singlet-Triplet energy gap, Hund’s rule violation, DFT, Exchange-Correlation
I Introduction
@@
The energy criterion for organic light-emitting diodes (OLEDs) to operate via thermally activated delayed fluorescence (TADF) is that the energy of the first excited singlet state (S1) must be no more than $0.1$ eV higher than that of the triplet state (T1Uoyama et al. (2012); Dos Santos et al. (2024). This condition facilitates thermally-aided reverse intersystem crossing (RISC), enabling the theoretical complete transfer of the T1 population to the emissive state, S1 Chen, Kim, and Brédas (2018). A derivative of the triangular molecule, heptazine (aka heptaazaphenalene, 7AP), with a central nitrogen (N) atom in an anti-aromatic 12-annulene framework, showed delayed fluorescence without thermal assistance, indicating a negative S1–T1 gap (STG) Aizawa et al. (2022). This mechanism paves the way for designing OLEDs that leverage exothermic delayed fluorescence from inverted singlet and triplet excited states (DFIST). Direct spectroscopic evidence of a negative STG ($-0.047\pm 0.007$ eV) was obtained for pentaazaphenalene (5AP) using anion photoelectron spectroscopy and fluorescence measurements Wilson et al. (2024). Negative STG was also observed in the transient photoluminescence spectrum of dialkylamine-substituted 5AP Kusakabe et al. (2024). Computational studies have further suggested that azaphenalenes (APs) with other substitution patterns, as well as the boron (B) analog of cyclazine (1AP), boraphenalene (1BP), also exhibit negative STGs Won, Nakayama, and Aizawa (2023); Li et al. (2022); Tučková et al. (2022); Dreuw and Hoffmann (2023); Ghosh and Bhattacharyya (2022); Ricci et al. (2021); Sancho-Garcia et al. (2022); Loos, Lipparini, and Jacquemin (2023). Non-alternant hydrocarbons and their substituted analogs also show the potential for negative STGs Garner, Blaskovits, and Corminboeuf (2023); Terence Blaskovits, Garner, and Corminboeuf (2023); Garner, Blaskovits, and Corminboeuf (2024); Blaskovits, Corminboeuf, and Garner (2024a); Nigam et al. (2024). In particular, substituted analogs of the non-fused bicyclic hydrocarbon have demonstrated negative STGs, which is attributed to through-bond charge-transfer states Blaskovits, Corminboeuf, and Garner (2024b).
@@
A comprehensive search across $12,880$ structurally diverse small molecules revealed no exceptions to Hund’s rule, indicating that achieving the electronic structure criteria for ${\rm STG}<0$ requires nontrivial electronic structure and molecular geometry Majumdar and Ramakrishnan (2024). Moreover, the typical magnitudes of negative STGs reported Loos, Lipparini, and Jacquemin (2023) based on first-principles calculations are smaller than $0.3$ eV, which is similar to the average error associated with popular density functional theory approximations (DFAs). Consequently, due to the rarity of molecules with negative STGs and the limitations of current quantum chemical approximations, accurately identifying molecules exhibiting ${\rm STG}<0$ via high-throughput screening, with minimal false-positive (i.e. actual STG is $>0$ but predicted to be $<0$) and false-negative (i.e. actual STG is $<0$ but predicted to be $>0$) predictions, remains challenging. The linear-response time-dependent density functional theory (LR-TD-DFT) formalism provides qualitative accuracy—i.e., gets the sign of the STG for DFIST candidates correct—only within the double-hybrid (dh) DFT framework containing many-body correlation via the second-order perturbation theory (MP2). With the right combination of ingredients, dh-DFT can reliably predict a wide range of molecular properties as benchmarked against higher-level methods or experimental information Tarnopolsky et al. (2008); Goerigk and Grimme (2011, 2014, 2011). Nevertheless, selecting suitable exchange-correlation (XC) functionals within the dh-DFT framework remains challenging, especially for novel problems lacking sufficient reference data.
@@
In this study, we revisit the basic mathematics of data fitting to show how to exploit the bias-variance tradeoff in model building to correct systematic errors in a model which is highly precise (i.e. low-variance) but has a low accuracy (i.e. high-bias). By exploiting the freedom of parameterization within the dh-DFT formalism, we propose to apply this strategy to model DFIST molecules with ${\rm STG}<0$. To this end, we evaluate various computational methods for predicting twelve STGs in triangular molecules, benchmarking them against previously reported theoretical best estimates (TBEs) based on composite models Loos, Lipparini, and Jacquemin (2023). We examine dh-DFAs exhibiting very small variance for STG predictions and reparametrize them to correct for their bias due to systematic errors. For this purpose, we also identify optimal low-bias dh-DFA parameterizations to use as internal references for correcting the predictions of the high-bias, low-variance formalisms. The utility of this internally referenced scaling approach is demonstrated by its ability to predict STGs with low bias and variance, resulting in excellent overall accuracy and precision. We finally discuss the merits and drawbacks of this approach for its further applications.
II Dataset and Computational Methods
We consider twelve systems (see Figure 1) using equilibrium geometries from Ref. Loos, Lipparini, and Jacquemin, 2023, determined at the coupled-cluster with singles, doubles, and perturbative triples (CCSD(T)) level, employing the frozen-core approximation and the cc-pVTZ basis set. Using these geometries, we calculated the electronic excited state energy levels with various methods in a single-point fashion. In addition, we collected TBEs of the S1 and T1 energies of the systems shown in Figure 1 from Ref. Loos, Lipparini, and Jacquemin, 2023 to serve as reference values for benchmarking.
TBE of STGs reported in Ref. Loos, Lipparini, and Jacquemin, 2023 are based on the composite procedure Loos, Lipparini, and Jacquemin (2023); Loos and Jacquemin (2021), where the S1 energy is extrapolated as CC3/aug-cc-pVTZ + [CCSDT/6-31+G(d) – CC3/6-31+G(d)], and the T1 energy is extrapolated as CCSD/aug-cc-pVTZ + [CC3/aug-cc-pVDZ – CCSD/aug-cc-pVDZ]. In other words, TBE of S1 energies were computed at the CC3/aug-cc-pVTZ level with a post-CC3 correction via EOM-CCSDT (using the 6-31+G(d) basis set). In contrast, the T1 energies were obtained at the EOM-CCSD/aug-cc-pVTZ level with a post-CCSD correction via CC3 using the aug-cc-pVDZ basis set Loos, Lipparini, and Jacquemin (2023). Although the TBE approach treats S1 and T1 energies using different protocols, both are estimated at the currently available higher levels of theory.
@@ II Dataset and Computational Methods
Figure 1: Nitrogen- and boron-centered [12]-annulene systems reported in Ref. Loos, Lipparini, and Jacquemin, 2023.
@@
Laplace-transformed, density-fitted, local versions of correlated methods, such as CC2 Freundorfer et al. (2010) and ADC(2) Schütz (2015), offer favorable speedups for modeling the electronic excited states of medium-sized organic molecules. We denote these methods as L-CC2 and L-ADC(2) and performed calculations using the cc-pVDZ, aug-cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets, along with the corresponding JKFIT and MP2FIT auxiliary basis sets using Molpro (version 2015.1) Werner et al. (2015). L-CC2 and L-ADC(2) are the same methods denoted as LT-DF-CC2 and LT-DF-ADC(2) elsewhere. L-CC2 has demonstrated excellent accuracy in modeling the excited states of BODIPY derivatives Feldt and Brown (2021); Momeni and Brown (2016). L-CC2 has been shown to deviate from the canonical variant CC2 by $\leq 0.05$ eV, irrespective of the character of the excited states, including that of charge-transfer states for which CC2’s predictions can be erroneous Kats, Korona, and Schütz (2006). The systems studied in this work do not feature charge-transfer characteristics, which is why CC2 predicts accurate STGs in agreement with TBEs reported in Ref. Loos, Lipparini, and Jacquemin, 2023.
For comparison, we collected S1 and T1 energies obtained using the linear-response (LR) coupled-cluster methods, EOM-CCSD and CC3, with the aug-cc-pVDZ basis set from Ref. Loos, Lipparini, and Jacquemin, 2023. Additionally, we performed ADC(2) calculations using the cc-pVDZ, cc-pVTZ, aug-cc-pVDZ, and aug-cc-pVTZ basis sets with the resolution-of-the-identity (RI) approximation Vahtras, Almlöf, and Feyereisen (1993); Kendall and Früchtl (1997) as implemented in QChem (version 6.0.2) Krylov and Gill (2013). As the excited state energy eigenvalues are typically calculated with iterative eigensolvers, at L-CC(2), L-ADC(2), and ADC(2) levels, we calculated the lowest three singlet and the lowest three triplet roots for sufficient variational convergence of the S1 and T1 states.
@@
We performed LR-TD-DFT calculations using the Tamm–Dancoff approximation (TDA). It is known that TDA may not predict oscillator strengths as accurately as the DFT version of random-phase approximation (RPA). However, RPA predictions of triplet energies are sensitive to triplet stability Rangel et al. (2017). Therefore, we prefer TDA for automated high-throughput calculations of excited-state energies. These energies may be combined with RPA-based oscillator strengths for more detailed analysis. Furthermore, previous work has demonstrated that its STG predictions for azaphenalenes are more accurate than those predicted using RPA Majumdar and Ramakrishnan (2024). The performance of various DFAs for predicting STGs in azaphenalenes has been discussed in prior studies Bhattacharyya (2021); Curtis et al. (2023); Kunze et al. (2024). Within the TDA formalism, we calculated the lowest six singlet and the lowest six triplet roots.
@@
This study explores several DFAs spanning all levels of the Jacob’s ladder Perdew and Schmidt (2001), namely, local density approximation (LDA): VWN5 Wilk, Nusair, and Vosko (1981) and PWLDA; generalized gradient approximation (GGA): BP86, BLYP, PW91 Perdew and Wang (1992); Burke, Perdew, and Wang (1998), PBE Perdew, Burke, and Ernzerhof (1996), mPWPW, and mPWLYP; hybrid GGA: B3LYP Stephens et al. (1994), PBE0 Adamo and Barone (1999), X3LYP Su, Xu, and Goddard (2004), mPW1PW, mPW1LYP, and BHandHLYP Becke (1993); meta-GGA: TPSS Perdew et al. (2009), TPSSh Perdew et al. (2009), TPSS0 Perdew et al. (2009), and M06L Zhao and Truhlar (2006) hybrid meta-GGA: M062X Zhao and Truhlar (2008); range-separated hybrids: $\omega$B97, Chai and Head-Gordon (2008) $\omega$B97X, Chai and Head-Gordon (2008)$\omega$B97X-D3, CAM-B3LYP Yanai, Tew, and Handy (2004), LC-BLYP Tawada et al. (2004), and LC-PBE Iikura et al. (2001); double hybrid: $\omega$B97X-2 Chai and Head-Gordon (2009), B2PLYP Grimme (2006), mPW2PLYP Schwabe and Grimme (2006), B2GP-PLYP Karton et al. (2008), PWPB95 Goerigk and Grimme (2011), PBE-QIDH Brémond et al. (2014), PBE0-DH Brémond and Adamo (2011), $\omega$B2PLYP Casanova-Páez, Dardis, and Goerigk (2019), $\omega$B2GP-PLYP Casanova-Páez, Dardis, and Goerigk (2019), $\omega$B88PP86 Casanova-Páez and Goerigk (2021a), $\omega$PBEPP86 Casanova-Páez and Goerigk (2021a), SCS/SOS-B2PLYP21 Casanova-Páez and Goerigk (2021a), SCS-PBE-QIDH Casanova-Páez and Goerigk (2021a), SOS-PBE-QIDH Casanova-Páez and Goerigk (2021a), SCS-B2GP-PLYP21 Casanova-Páez and Goerigk (2021a), SOS-B2GP-PLYP21 Casanova-Páez and Goerigk (2021a), SCS-$\omega$B2GP-PLYP21 Casanova-Páez and Goerigk (2021a), SOS-$\omega$B2GP-PLYP21 Casanova-Páez and Goerigk (2021a), SCS-RSX-QIDH Casanova-Páez and Goerigk (2021a), SOS-RSX-QIDH Casanova-Páez and Goerigk (2021a), SCS-$\omega$B88PP86 Casanova-Páez and Goerigk (2021a), SOS-$\omega$B88PP86 Casanova-Páez and Goerigk (2021a), SCS-$\omega$PBEPP86 Casanova-Páez and Goerigk (2021a), SOS-$\omega$PBEPP86 Casanova-Páez and Goerigk (2021a), DSD-BLYP Yu (2014), DSD-PBEP86 Yu (2014), DSD-PBEB95 Yu (2014), RSX-QIDH Bremond et al. (2018), and RSX-0H Brémond et al. (2019). Here, SCS and SOS denote spin-component-scaled and spin-opposite-scaled, respectively. Such corrections have generally been shown to improve the MP2 correlation component in dh-DFAs. All DFT calculations were performed using Orca (version 6.0.0) Neese (2012, 2018).
III Bias Correction for Low-Variance Predictors
Generally, when developing a mathematical model, such as deductive models based on quantum mechanical approximations or inductive models based on data fitting or machine learning, a desired aspect is that the resulting mathematical model is both precise (with low variance) and accurate (with low bias or systematic error). However, due to the trade-offs involved in incorporating physical effects while keeping the model computationally tractable, computational chemistry models, particularly DFAs, often exhibit a varying degree of systematic errors across different formulations, such as local density approximation (LDA), generalized gradient approximation (GGA), hybrid-GGA, and so on. This aspect has been highlighted for numerous properties, such as binding energies across alkanes Wodrich, Corminboeuf, and Schleyer (2006), spin-splitting energies Hughes and Friesner (2011), chemical reaction energies Simm and Reiher (2016), to name a few instances. For atomization or formation energies, DFT-based predictions are often corrected for systematic errors via “dressed atom corrections”, facilitating decreasing across DFAs while climbing higher on the “Jacob’s ladder of DFAs” Winget and Clark (2004); Grimme (2006); Das, Chakraborty, and Ramakrishnan (2021).
The main goal of this study is to derive formal relations that enable one to correct a low-variance, high-bias model (i.e. precise but inaccurate) by correcting it through linear regression using a high-variance, low-bias model as an internal reference. We begin by examining the conditions under which such an internally referenced bias correction is effective. To the best of our knowledge, the details of this procedure have not been reported elsewhere. However, given the simplicity of the technique proposed here, it could be included in a plethora of calibration techniques used in data-fitting Wohland, Maiti, and Macháň (2020).
III.1 Definitions